# On semipositone problems over ${\mathbb{R}}^{N}$ for the fractional $p$-Laplace operator

### v3

@@ Abstract.
For $N\geq 1,s\in(0,1)$, and $p\in(1,N/s)$ we find a positive solution to the following class of semipositone problems associated with the fractional $p$-Laplace operator:
@@
where $g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ is a positive function, $a>0$ is a parameter and $f_{a}\in{\mathcal{C}}(\mathbb{R})$ is defined as $f_{a}(t)=f(t)-a$ for $t\geq 0$, $f_{a}(t)=-a(t+1)$ for $t\in[-1,0]$, and $f_{a}(t)=0$ for $t\leq-1$, where $f\in{\mathcal{C}}(\mathbb{R}^{+})$ satisfies $f(0)=0$ with subcritical and Ambrosetti-Rabinowitz type growth. Depending on the range of $a$, we obtain the existence of a mountain pass solution to (SP) in $\mathcal{D}^{s,p}(\mathbb{R}^{N})$. Then, we prove mountain pass solutions are uniformly bounded with respect to $a$, over $L^{r}(\mathbb{R}^{N})$ for every $r\in[Np/N-sp,\infty]$. In addition, if $p>2N/N+2s$, we establish that (SP) admits a non-negative mountain pass solution for each $a$ near zero. Finally, under the assumption $g(x)\leq B/|x|^{\beta(p-1)+sp}$ for $B>0,x\neq 0$, and $\beta\in(N-sp/p-1,N/p-1)$, we derive an explicit positive radial subsolution to (SP) and show that the non-negative solution is positive a.e. in $\mathbb{R}^{N}$.
@@ 1. Introduction
$f_{a}(t)=\begin{cases}f(t)-a&\text{ if }t\geq 0,\\ -a(t+1)&\text{ if }t\in[-1,0],\\ 0&\text{ if }t\leq-1,\end{cases}\text{ where }f\in{\mathcal{C}}([0,\infty))\text{ with }f(0)=0.$ (1.1)
Also, $f$ satisfies the following growth assumptions:
@@
${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}:=\overline{C_{c}^{\infty}({\mathbb{R}}^{N})}^{\left\lVert\cdot\right\rVert_{s,p}},\text{ where }\|u\|_{s,p}:=\bigg{(}\iint_{{\mathbb{R}}^{2N}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{N+sp}}\ {\mathrm{d}x\mathrm{d}y}\bigg{)}^{\frac{1}{p}}.$
@@
The study of semipositone problems began from the work [13] of Brown and Shivaji while studying the bifurcation theory for the perturbed problem $-\Delta u=\lambda(u-u^{3})-\epsilon$ in $\Omega$ and $u>0$ in $\Omega$, where $\lambda,\epsilon>0$ and $\Omega$ is a bounded domain. Subsequently, numerous authors have investigated the existence and the qualitative aspects of positive solutions to local semipositone problems across various domains. For bounded domain $\Omega$, relevant studies can be found in [21, 18, 15, 17, 19, 23, 3, 14, 35, 28, 22, 38, 33]. In the context where $\Omega$ is the exterior of a bounded domain, we refer [20, 36, 32, 31, 24] and the references therein. In [2], Alves et al. first studied the semipositone problem within the entire domain ${\mathbb{R}}^{N}$, as described by the equation $-\Delta u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N};N\geq 3$ with $f(0)=0$ and $a>0$. The function $f\in\mathcal{C}({\mathbb{R}}^{+})$ is locally Lipschitz, has subcritical and the Ambrosetti-Rabinowitz (A-R) type growth. The weight $g$ is positive and bounded by a radial function $P\in{\mathcal{C}}({\mathbb{R}}^{+})$, where $P$ satisfies $(a)\int_{{\mathbb{R}}^{N}}|x|^{2-N}P(|x|)\,{\rm d}x<\infty$, $(b)$ $P(\lvert\cdot\rvert)\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$, and $(c)$ $|x|^{N-2}\int_{{\mathbb{R}}^{N}}P(|y|)/|x-y|^{N-2}\,{\mathrm{d}y}\leq C$, for all $x\in{\mathbb{R}}^{N}\setminus\{0\}$ and for some constant $C>0$. For $p\in(1,N)$, in [1], Santos et al. investigated the nonlinear variant $-\Delta_{p}u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N}$, where $f(0)=0,f\in{\mathcal{C}}({\mathbb{R}}^{+})$ exhibits subcritical with A-R type growth and the weight $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ satisfies $g(x)<B|x|^{-\vartheta}$ for $x\neq 0$, with $\vartheta>N$ and $B>0$. Meanwhile, the study in [6] focused on $\Delta^{2}u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N};N\geq 5$ where $f(0)=0$, $f\in{\mathcal{C}}({\mathbb{R}}^{+})$ satisfies weaker A-R type growth.
@@
Few articles are available in the literature dealing with non-local semipositone problems. In this direction, for $N>2s$ and $\Omega$ bounded, the authors in [25] studied $(-\Delta)^{s}u=\lambda(u^{q}-1)+\mu u^{r}$ in $\Omega$; $u=0$ in ${\mathbb{R}}^{N}\setminus\Omega$, where $\lambda,\mu>0$, $q\in(0,1)$ and $r\in(1,\frac{N+2s}{N-2s})$. Under certain lower bound of $\lambda$, they constructed a positive subsolution when $\mu=0$ and showed that there exists at least one positive solution when $0<\mu<\mu_{\lambda}$. For $N>p\geq 2$ and $\Omega$ bounded, the authors in [34] proved that $(-\Delta)_{p}^{s}u=\lambda f(u)$ in $\Omega$; $u=0$ in ${\mathbb{R}}^{N}\setminus\Omega$ admits positive solution, provided $\lambda>0$ is small. They used regularity of weak solution up to the boundary of $\Omega$ and Hopf’s Lemma for $(-\Delta)_{p}^{s}$. Recently, in [5], the author studies $(-\Delta)^{s}u=g(x)\left(f(u)-a\right)$ in $\mathbb{R}^{N}$ with $f(0)=0,f\in{\mathcal{C}}({\mathbb{R}}^{+})$ is locally Lipschitz, satisfies subcritical and weaker A-R type growth. Whereas, $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ satisfies $|x|^{N-2s}\int_{{\mathbb{R}}^{N}}g(y)/|x-y|^{N-2s}\,{\mathrm{d}y}\leq C(g),x\in{\mathbb{R}}^{N}\setminus\{0\},C(g)>0$. The existence of a positive solution is obtained employing similar techniques as in [2].
In this paper, we aim to study (SP), a non-local analogue of [1]. To our knowledge, non-linear non-local semipositone problems on the whole of $\mathbb{R}^{N}$ have not been addressed in the literature. The weight function $g$ falls within both $L^{1}(\mathbb{R}^{N})$ and $L^{\infty}(\mathbb{R}^{N})$ spaces, adheres to the bounds specified in (1.4). Meanwhile, the function $f$ meets subcritical and A-R type growth, as outlined in ($\bf f{1}$) and ($\bf f{2}$). Depending on the parameter $a$, our principal goal is to establish the existence of a positive solution to (SP). The nonsmooth variational technique as in [1] is not readily adapted for $s\in(0,1)$ due to the regularity constraints associated with the critical points of the non-smooth energy functional of (SP). This leads us to follow a different approach from [1]. We consider an energy functional associated with (SP), which has a ${\mathcal{C}}^{1}$ variational structure (see (2.1)). We say a weak solution $u$ is a mountain pass solution of (SP) if it is a critical point of the energy functional, which satisfies the mountain pass geometry and the Palais-Smale condition (see Proposition 2.4 and Proposition 2.6). The following theorem combines our main results.
@@
To show the positivity of the solution $u_{a}$, we first show that $\{u_{a}\}$ uniformly converges to a positive function in ${\mathcal{C}}({\mathbb{R}}^{N})$ as $a\rightarrow 0$ (see Proposition 3.5, Proposition 3.6, and Proposition 3.8). As $a\rightarrow 0$, we obtain an explicit positive subsolution of (SP) on the exterior of a ball in ${\mathbb{R}}^{N}$ following the approach in [11, Lemma 3.4]. Subsequently, we obtain the positivity of $u_{a}$ using the comparison principle [11, Theorem 2.7]. The strategy to prove Theorem 1.1-((c) and (d)) is different from [5, 2], where the Riesz potential for a linear operator plays a major role.
@@ 2. Functional frameworks for the problem
(iv) For $p\in[1,\infty]$, $L^{p}({\mathbb{R}}^{N})$ norm of a function $u$ is denoted as $\left\lVert u\right\rVert_{p}$.
@@
Moreover, the critical point of $I_{a}$ corresponds to the solution of (SP). Before discussing further properties of $K_{a},I_{a}$, we identify the upper and lower bounds of $f,f_{a},$ and their primitives. Clearly, the function $f_{a}\in{\mathcal{C}}({\mathbb{R}})$ and its primitive $F_{a}$ is defined as $F_{a}(t)=F(t)-at$ for $t\geq 0$, $F_{a}(t)=-\frac{at^{2}}{2}-at$ for $t\in[-1,0]$ and $F_{a}(t)=\frac{a}{2}$ for $t\leq-1$. From ($\bf f{1}$),
@@
Again using the subcritical growth on $f$, $f(t)\leq C(f)t^{\gamma-1}$ for $t>t_{2}$. The continuity of $f$ infers that $f(t)\leq C$ on $[0,t_{2}]$. Hence for $a\in(0,\tilde{a})$, we obtain
@@
(ii) Assume that $u_{n}\rightharpoonup u$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. We need to prove $u_{n}\rightarrow u$ in $L^{q}({\mathbb{R}}^{N},|g|)$. The space ${\mathcal{C}}_{c}({\mathbb{R}}^{N})$ is dense in $L^{\alpha}({\mathbb{R}}^{N})$ and hence for every $\epsilon>0$ we take $g_{\epsilon}\in{\mathcal{C}}_{c}({\mathbb{R}}^{N})$ with $K:=\mathrm{supp}(g)$ such that
@@
The above constant $M$ is the upper bound of $g_{\epsilon}$ on the compact set $K$. Moreover, the embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{q}_{\text{loc}}({\mathbb{R}}^{N})$ is compact. Therefore, there exists $n_{1}\in{\mathbb{N}}$ such that
@@
From (2) and (2.7), we deduce the following
@@
where $L:=\sup\{\|u_{n}\|_{p_{s}^{*}}+\|u\|_{p_{s}^{*}}+\|u_{n}\|_{p_{s}^{*}}^{\gamma}+\|u\|_{p_{s}^{*}}^{\gamma}:n\in{\mathbb{N}}\}$. For $a\geq 0$, we write
@@
    There exists $v\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ with $\|v\|_{s,p}>\beta$ such that $I_{a}(v)<0$.
@@
$\int_{{\mathbb{R}}^{N}}g(x)|u_{n}-u||u_{n}|^{\gamma-1}\,{\rm d}x\leq\left\lVert g\right\rVert_{\infty}^{\frac{\gamma-1}{p^{*}_{s}}}\left\lVert u_{n}-u\right\rVert_{L^{\frac{p_{s}^{*}}{p_{s}^{*}-(\gamma-1)}}({\mathbb{R}}^{N},g)}\|u_{n}^{\gamma-1}\|_{\frac{p_{s}^{*}}{\gamma-1}}\rightarrow 0\text{ as }n\rightarrow\infty.$ (2.18)

### v4

@@ Abstract.
For $N\geq 1,s\in(0,1)$, and $p\in(1,\frac{N}{s})$ we find a positive solution to the following class of semipositone problems associated with the fractional $p$-Laplace operator:
@@
where $g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ is a positive function, $a>0$ is a parameter and $f_{a}\in{\mathcal{C}}(\mathbb{R})$ is defined as $f_{a}(t)=f(t)-a$ for $t\geq 0$, $f_{a}(t)=-a(t+1)$ for $t\in[-1,0]$, and $f_{a}(t)=0$ for $t\leq-1$, where $f$ is a non-negative continuous function on $[0,\infty)$ satisfies $f(0)=0$ with subcritical and Ambrosetti-Rabinowitz type growth. Depending on the range of $a$, we obtain the existence of a mountain pass solution to (SP) in $\mathcal{D}^{s,p}(\mathbb{R}^{N})$. Then, we prove mountain pass solutions are uniformly bounded with respect to $a$, over $L^{r}(\mathbb{R}^{N})$ for every $r\in\left[\frac{Np}{N-sp},\infty\right]$. In addition, if $p>\frac{2N}{N+2s}$, we establish that (SP) admits a non-negative mountain pass solution for each $a$ near zero. Finally, under the assumption $g(x)\leq\frac{B}{|x|^{\beta(p-1)+sp}}$ for $B>0,x\neq 0$, and $\beta\in\left(\frac{N-sp}{p-1},\frac{N}{p-1}\right)$, we derive an explicit positive radial subsolution to (SP) and show that the non-negative solution is positive a.e. in $\mathbb{R}^{N}$.
@@ 1. Introduction
$f_{a}(t)=\begin{cases}f(t)-a&\text{ if }t\geq 0,\\ -a(t+1)&\text{ if }t\in[-1,0],\\ 0&\text{ if }t\leq-1,\end{cases}$ (1.1)
where $f$ is a non-negative continuous function on $[0,\infty)$ with $f(0)=0$. Also, $f$ satisfies the following growth assumptions:
@@
${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}:=\overline{{\mathcal{C}}_{c}^{\infty}({\mathbb{R}}^{N})}^{\left\lVert\cdot\right\rVert_{s,p}},\text{ where }\|u\|_{s,p}:=\bigg{(}\iint_{{\mathbb{R}}^{2N}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{N+sp}}\ {\mathrm{d}x\mathrm{d}y}\bigg{)}^{\frac{1}{p}}.$
@@
The study of semipositone problems began from the work [13] of Brown and Shivaji while studying the bifurcation theory for the perturbed problem $-\Delta u=\lambda(u-u^{3})-\epsilon$ in $\Omega$ and $u>0$ in $\Omega$, where $\lambda,\epsilon>0$ and $\Omega$ is a bounded domain. Subsequently, numerous authors have investigated the existence and the qualitative aspects of positive solutions to local semipositone problems across various domains. For bounded domain $\Omega$, relevant studies can be found in [21, 18, 15, 17, 19, 23, 3, 14, 36, 28, 22, 39, 34]. In the context where $\Omega$ is the exterior of a bounded domain, we refer [20, 37, 33, 32, 24] and the references therein. In [2], Alves et al. first studied the semipositone problem within the entire domain ${\mathbb{R}}^{N}$, as described by the equation $-\Delta u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N};N\geq 3$ with $f(0)=0$ and $a>0$. The function $f\in\mathcal{C}({\mathbb{R}}^{+})$ is locally Lipschitz, has subcritical and the Ambrosetti-Rabinowitz (A-R) type growth. The weight $g$ is positive and bounded by a radial function $P\in{\mathcal{C}}({\mathbb{R}}^{+})$, where $P$ satisfies $(a)\int_{{\mathbb{R}}^{N}}|x|^{2-N}P(|x|)\,{\rm d}x<\infty$, $(b)$ $P(\lvert\cdot\rvert)\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$, and $(c)$ $|x|^{N-2}\int_{{\mathbb{R}}^{N}}P(|y|)|x-y|^{-N+2}\,{\mathrm{d}y}\leq C$, for all $x\in{\mathbb{R}}^{N}\setminus\{0\}$ and for some constant $C>0$. For $p\in(1,N)$, in [1], Santos et al. investigated the nonlinear variant $-\Delta_{p}u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N}$, where $f(0)=0,f\in{\mathcal{C}}({\mathbb{R}}^{+})$ exhibits subcritical with A-R type growth and the weight $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ satisfies $g(x)<B|x|^{-\vartheta}$ for $x\neq 0$, with $\vartheta>N$ and $B>0$. Meanwhile, the study in [6] focused on $\Delta^{2}u=g(x)(f(u)-a)$, $u>0$ in $\mathbb{R}^{N};N\geq 5$ where $f(0)=0$, $f\in{\mathcal{C}}({\mathbb{R}}^{+})$ satisfies weaker A-R type growth.
@@
Few articles are available in the literature dealing with non-local semipositone problems. In this direction, for $N>2s$ and $\Omega$ bounded, the authors in [25] studied $(-\Delta)^{s}u=\lambda(u^{q}-1)+\mu u^{r}$ in $\Omega$; $u=0$ in ${\mathbb{R}}^{N}\setminus\Omega$, where $\lambda,\mu>0$, $q\in(0,1)$ and $r\in(1,\frac{N+2s}{N-2s})$. Under certain lower bound of $\lambda$, they constructed a positive subsolution when $\mu=0$ and showed that there exists at least one positive solution when $0<\mu<\mu_{\lambda}$. For $N>p\geq 2$ and $\Omega$ bounded, the authors in [35] proved that $(-\Delta)_{p}^{s}u=\lambda f(u)$ in $\Omega$; $u=0$ in ${\mathbb{R}}^{N}\setminus\Omega$ admits positive solution, provided $\lambda>0$ is small. They used regularity of weak solution up to the boundary of $\Omega$ and Hopf’s Lemma for $(-\Delta)_{p}^{s}$. Recently, in [5], the author studies $(-\Delta)^{s}u=g(x)\left(f(u)-a\right)$ in $\mathbb{R}^{N}$ with $f(0)=0,f\in{\mathcal{C}}({\mathbb{R}}^{+})$ is locally Lipschitz, satisfies subcritical and weaker A-R type growth. Whereas, $g\in L^{1}({\mathbb{R}}^{N})\cap L^{\infty}({\mathbb{R}}^{N})$ satisfies $|x|^{N-2s}\int_{{\mathbb{R}}^{N}}g(y)|x-y|^{-N+2s}\,{\mathrm{d}y}\leq C(g),x\in{\mathbb{R}}^{N}\setminus\{0\},C(g)>0$. The existence of a positive solution is obtained employing similar techniques as in [2].
In this paper, we aim to study (SP), a non-local analogue of [1]. To our knowledge, non-linear non-local semipositone problems on the whole of $\mathbb{R}^{N}$ have not been addressed in the literature. The weight function $g$ falls within both $L^{1}(\mathbb{R}^{N})$ and $L^{\infty}(\mathbb{R}^{N})$ spaces, adheres to the bounds specified in (1.4). Meanwhile, the function $f$ meets subcritical and A-R type growth, as outlined in ($\bf f{1}$) and ($\bf f{2}$). Depending on the parameter $a$, our principal goal is to establish the existence of a positive solution to (SP). The nonsmooth variational technique as in [1] is not readily adapted for $s\in(0,1)$ due to the regularity constraints associated with the critical points of the non-smooth energy functional of (SP). This leads us to follow a different approach from [1]. We consider an energy functional associated with (SP), which has a ${\mathcal{C}}^{1}$ variational structure (see (2.1)). Applying the mountain pass theorem, we establish the existence of a mountain pass critical point for the energy functional, which corresponds to a mountain pass solution of (SP). The following theorem combines our main results.
@@
To show the positivity of the solution $u_{a}$, we first show that the sequence of solutions $\{u_{a}\}$ uniformly converges to a positive function in ${\mathcal{C}}({\mathbb{R}}^{N})$ as $a\rightarrow 0$ (see Proposition 3.5, Proposition 3.6, and Proposition 3.8). For $a$ near zero, we obtain an explicit positive subsolution of (SP) on the exterior of a ball in ${\mathbb{R}}^{N}$, following the approach in [11, Lemma 3.4]. Subsequently, we obtain the positivity of $u_{a}$ using the comparison principle [11, Theorem 2.7]. The strategy to prove Theorem 1.1-((c) and (d)) is different from [5, 2], where the Riesz potential for a linear operator plays a major role.
@@ 2. Functional frameworks for the problem
(iv) For $p\in[1,\infty]$, the $L^{p}({\mathbb{R}}^{N})$ norm of a function $u$ is denoted as $\left\lVert u\right\rVert_{p}$.
@@
Moreover, every critical point of $I_{a}$ corresponds to a solution of (SP). Before discussing further properties of $K_{a},I_{a}$, we identify the upper and lower bounds of $f,f_{a},$ and their primitives. Clearly, the function $f_{a}\in{\mathcal{C}}({\mathbb{R}})$ and its primitive $F_{a}$ is defined as $F_{a}(t)=F(t)-at$ for $t\geq 0$, $F_{a}(t)=-\frac{at^{2}}{2}-at$ for $t\in[-1,0]$ and $F_{a}(t)=\frac{a}{2}$ for $t\leq-1$. From ($\bf f{1}$),
@@
Again using the subcritical growth on $f$, $f(t)\leq C(f)t^{\gamma-1}$, for $t>t_{2}$, for some $t_{2}>0$. The continuity of $f$ infers that $f(t)\leq C$ on $[0,t_{2}]$. Hence for $a\in(0,\tilde{a})$, we obtain
@@
(ii) Assume that $u_{n}\rightharpoonup u$ in ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$. We need to prove $u_{n}\rightarrow u$ in $L^{q}({\mathbb{R}}^{N},|g|)$. The space ${\mathcal{C}}_{c}({\mathbb{R}}^{N})$ is dense in $L^{\alpha}({\mathbb{R}}^{N})$ and hence for every $\epsilon>0$ we take $g_{\epsilon}\in{\mathcal{C}}_{c}({\mathbb{R}}^{N})$ with $K:=\mathrm{supp}(g_{\epsilon})$ such that
@@
The above constant $M$ is the upper bound of $g_{\epsilon}$ on the compact set $K$. Moreover, the embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{q}_{\text{loc}}({\mathbb{R}}^{N})$ is compact. Therefore, there exists $n_{1}\in{\mathbb{N}}$ such that up to a subsequence
@@
From (2) and (2.7), we deduce that
@@
where $L:=\sup\{\|u_{n}\|_{p_{s}^{*}}+\|u\|_{p_{s}^{*}}+\|u_{n}\|_{p_{s}^{*}}^{\gamma}+\|u\|_{p_{s}^{*}}^{\gamma}:n\in{\mathbb{N}}\}$. Since the sequence $\{\|u_{n}\|_{s,p}\}$ is bounded in $\mathbb{R}^{+}$, using the continuous embedding ${{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}\hookrightarrow L^{p_{s}^{*}}({\mathbb{R}}^{N})$ we see that $\{\|u_{n}\|_{p_{s}^{*}}\}$ is also bounded in $\mathbb{R}^{+}$. Therefore, $L<\infty$. For $a\geq 0$, we write
@@
    There exists $v\in{{\mathcal{D}}^{s,p}({\mathbb{R}}^{N})}$ with $\|v\|_{s,p}>\beta$ such that $I_{a}(v)<0$, for all $a>0$.
@@
$\int_{{\mathbb{R}}^{N}}g(x)|u_{n}-u||u_{n}|^{\gamma-1}\,{\rm d}x\leq\left\lVert g\right\rVert_{\infty}^{\frac{\gamma-1}{p^{*}_{s}}}\left\lVert u_{n}-u\right\rVert_{L^{\frac{p_{s}^{*}}{p_{s}^{*}-(\gamma-1)}}({\mathbb{R}}^{N},g)}\||u_{n}|^{\gamma-1}\|_{\frac{p_{s}^{*}}{\gamma-1}}\rightarrow 0\text{ as}n\rightarrow\infty.$ (2.18)